OpenStudy (anonymous):

Find all values of x in the interval [0, 2π] that satisfy the equation. (Enter your answers as a comma-separated list.) 18 + 9 cos(2x) = 27 cos(x)

OpenStudy (xapproachesinfinity):

hint subtract 27cosx from both side

OpenStudy (xapproachesinfinity):

i bring to a point where it is easy to do 9cos2x-27cosx+18=0 cos2x-3cosx+2=0

OpenStudy (xapproachesinfinity):

cos2x=2cos^2x-1

OpenStudy (xapproachesinfinity):

so we have 2cos^2x-3cosx+1=0

OpenStudy (xapproachesinfinity):

surely you can do an equation like this 2x^2-3x+1=0

OpenStudy (xapproachesinfinity):

can you?

OpenStudy (anonymous):

Aha, Thank you very much for replying. Let me try it and get back to you

OpenStudy (xapproachesinfinity):

ok

OpenStudy (anonymous):

I get it until 2cos^2x-3cosx+1=0

OpenStudy (xapproachesinfinity):

alright now we just think of that as 2r^2-3r+1=0 for now we forgot that cosx like we replaced with r

OpenStudy (anonymous):

okay...

OpenStudy (xapproachesinfinity):

2r^2-3r+1=0 is just quadratic equation

OpenStudy (anonymous):

ohh I see...

OpenStudy (anonymous):

okay so r = 1/2 or 1 yeah?

OpenStudy (xapproachesinfinity):

you solve for r just as you usually do -factoring -quadratic formula -completing the square i don't know what other stuff you can use lol

OpenStudy (xapproachesinfinity):

yes seems good

OpenStudy (anonymous):

lol...okay, so replacing r with cos will result in...?

OpenStudy (xapproachesinfinity):

if i did factoring we get (2r-1)(r-1)=0 which agrees with what you got r=1/2 and r=1 we replace r by cosx cosx=1/2 and cosx=1

OpenStudy (xapproachesinfinity):

now you think of all possible angles that give 1/2 and 1 from 0 to 2pi

OpenStudy (xapproachesinfinity):

forr what angles do we have cos =1/2?

OpenStudy (anonymous):

60

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

and cos 0 = 1

OpenStudy (xapproachesinfinity):

not onnly 0 though

OpenStudy (xapproachesinfinity):

360 too

OpenStudy (anonymous):

in radians that will be pi/3...I mean for 60...yeah?

OpenStudy (xapproachesinfinity):

yeah

OpenStudy (anonymous):

yeah...360 too

OpenStudy (xapproachesinfinity):

360 or 2pi in radian

OpenStudy (anonymous):

that will be 2pi

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

cos 300...5pi/3

OpenStudy (xapproachesinfinity):

then x=0,pi/3,2pi

OpenStudy (anonymous):

that gives 1/2 yeah?

OpenStudy (xapproachesinfinity):

we only need values from 0 to 2pi and we stop we don;t need other rotations

OpenStudy (anonymous):

yeah...but 5pi/3 is actually smaller than 2pi

OpenStudy (xapproachesinfinity):

i was giving you how to put it as a complete answer with commas as they asked x=0,pi/3, 2pi

OpenStudy (xapproachesinfinity):

oh yeah right that has cos =1/2 too correct i missed that one

OpenStudy (xapproachesinfinity):

good catch

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

THANK YOU SO SO SO MUCH...I AM REALLY GREATFUL!!!! :-)

OpenStudy (xapproachesinfinity):

no problem

OpenStudy (anonymous):

:-)

OpenStudy (xapproachesinfinity):

until next time :) thanks for the testimonial

OpenStudy (anonymous):

Sure!!! :-D

OpenStudy (anonymous):

Please fan me so I can send you messages...If you don't mind though! @xapproachesinfinity

OpenStudy (xapproachesinfinity):

sure :)

OpenStudy (anonymous):

Thank You :-)...@xapproachesinfinity